# Roberto Dale,22footnotetext: Corresponding author.    Jesús M. Gandía    Juan Antonio Morales-Lladosa    and Ramon Lapiedra

### v1

@@ Abstract
In a recent paper [R. Dale. R. Lapiedra, and J. A. Morales-Lladosa, Phys. Rev. D 107, 023506 (2023)] a cosmic-like Clauser-Horne-Shimony-Holt (CHSH) inequality was proved for the temperature fluctuations $\delta T$ of the perturbed Cosmic Microwave Background (CMB), assuming local realism. This inequality can be tested from observational data. In fact, no violation for it has been found from the CMB data provided by the COBE satellite in different sky directions. This is a result which, being negative, is not conclusive in relation to the possible violation of the CHSH inequality in a cosmological context. So, the preceding analysis needs to be extended by using more precise current CMB data, like those provided by WMAP and Planck satellites. This is a task of considerable importance but, technically, much more involved in data treatments. In this work, assuming again local realism behind measurements and observational data, this extension is accomplished. The result is that a sound violation of the CHSH inequality is found, which would mean the failure of the assumed local realism in accordance with the quantum origin of the primordial temperature fluctuations in the framework of standard inflation and the orthodox interpretation of the Quantum Mechanics.
@@ 1 Introduction
Hence, it is usually assumed that a primordial quantum distribution of cosmic fluctuations dominated a very early era of the Universe but that CMB observational data respond to a classical statistic treatment obeying a stochastic distribution [22]. Diverse proposals exist to explain how such a quantum to classical transition takes place [22, 23, 24, 25], which are not free of criticisms [26]. Moreover, it is important to assess the degree of quantumness, if any, that a scalar inflationary field and the observed CMB anisotropies could encode [27, 28]. The transition from the firstborn quantum era to a classical o semi-classical one is an important topic which continues being addressed today, both from the theoretical and observational points of views [8, 29]. Thus, the proposed theoretical models go forward to an inasmuch more interesting, but very “entangled” issue in Cosmology which claims for an experimental resolution.
@@
This work does not include new theoretical insight or discussion about the transition from a quantum phase to a radiation dominant era, nor about how the primordial fluctuations were imprinted as CMB fluctuations, and its posterior evolution in an expanding universe. On the contrary, to face the issue we concentrate on the possibility to establish the true nature, if any, of the CMB anisotropies from their own observation. We propose a new bet to solve this “entangled” cosmic situation: to cut the “Gordian knot” with the sword of experiment.
@@
The present paper focuses on a variant of Bell’s inequality [30], known as the CHSH inequality (named after its authors, [31]), whose proof has been revisited in [32]. This inequality has been suitably mimicked and adapted to deal with CMB anisotropies data; then it was named “cosmic CHSH inequality”. In Ref. [10] no violation of this inequality was reported from the analysis of the COBE mission data [33]. But a concluding result about the inequality fulfillment deserves fuller treatment and demands a great amount of data like those provided by the WMAP and Planck measurement results [34, 35]. Now, in both cases, we find a sound violation of the Bell inequalities, which could be interpreted as a failure of local realism. The realism assumption posits that measurements can reveal an objective physical property intrinsic to the CMB, which exists prior to and independently of any observation. The locality assumption asserts that measurements conducted in two spacelike-separated regions at the decoupling time cannot exert any physical influence on each other at any subsequent cosmic instant.
@@
Laboratory experiments concerning the violation of Bell inequalities (see [11, 36] and references therein) show favour to Born’s interpretation of quantum mechanics over the de Broglie-Bohm interpretation. Nevertheless it seems to us that the controversy between both quantum interpretations continues open, particularly in Cosmology [37, 38]. Thus, the present paper could serve as a cosmic referee, potentially motivating the exploration of loopholes [39, 40, 41, 42, 43, 44] and alternative classical interpretations [45, 46], or ruling them out [47, 48], thereby providing new insights into the longstanding debate. This revision improves readability, precision, and fluency while maintaining a formal tone appropriate for scientific writing.
@@
The paper is organised as follows. First, Section 2 examines the fundamental hypothesis of local realism, which underpins the proof of Bell’s theorem, and reconciles it with measurements of the CMB. This section also revisits the “cosmic CHSH inequality”. Section 3 addresses the calculation of probabilities, expected values, and statistical methods applied to these cosmic inequalities. The computational procedures are summarised in Section 4. Subsequently, Sections 5 and 6 analyse observational data from the WMAP and Planck missions, demonstrating violations of the cosmic CHSH inequality in both cases. Section 7 provides a concise summary of the results obtained using degraded maps, including the confirmation of these findings with the original maps. Finally, Section 8 discusses the assumptions considered and the physical conclusions derived. Supplementary computations and technical results are presented in Appendix A.
@@ 2.1 Local realism assumption
Let $A$ and $B$ denote events in a spacelike configuration at the decoupling time, and let $\boldsymbol{\hat{x}}$ and $\boldsymbol{\hat{y}}$ be unit vectors associated with past-oriented light-like directions pointing towards $A$ and $B$, respectively. Following Ref. [10], we consider the dichotomised values of $\delta T(\boldsymbol{\hat{x}})$, defined by the function $F$, such that $F(\boldsymbol{\hat{x}})=+1$ (or $-1$) if $\delta T(\boldsymbol{\hat{x}})$ is positive (or negative). A similar definition applies for the sky direction $\boldsymbol{\hat{y}}$. To derive a cosmic CHSH inequality, the hypothesis of local realism is assumed.
By definition, realism implies the existence of hidden variables, $\lambda$, predating the measurements, such that the dichotomised CMB anisotropy values (±$\pm 1$) are well-defined functions of $\lambda$ and the corresponding measurement directions. Explicitly,
@@
After each pair of measurements is completed, a new run is initiated, without requiring the subsequent pair of measurements to be conducted in the same initial state as the previous one. In a laboratory-based CHSH-like experiment, the system is actively prepared in the desired initial state before each new run. In the cosmic case, however, the invariance of the initial conditions prior to any measurement is not a crucial factor. This is because successive CMB measurement pairs are taken along randomly selected sky directions that maintain a constant angular separation. The only requirement is the assignment of a value, either $+1$ or $-1$, to each measurement.
For each run, $\boldsymbol{\hat{x}}$ randomly assumes one of two possible values, $\boldsymbol{\hat{a}}$ or $\boldsymbol{\hat{a}^{\prime}}$. Similarly and independently, $\boldsymbol{\hat{y}}$ takes one of two values, $\boldsymbol{\hat{b}}$ or $\boldsymbol{\hat{b}^{\prime}}$. We then consider the expected values $\left\langle\delta T_{A}\delta T_{B}\right\rangle_{(\boldsymbol{\hat{x}\hat{y}})}$ for the four possible cases: $(\boldsymbol{\hat{a}\hat{b}})$, $(\boldsymbol{\hat{a}\hat{b}^{\prime}})$, $(\boldsymbol{\hat{a}^{\prime}\hat{b}})$, and $(\boldsymbol{\hat{a}^{\prime}\hat{b}^{\prime}})$. The locality assumption reflects the spacelike separation of the two measurement events during each run.
@@ 2.2 Statment of the cosmic CHSH inequality
Let us summarize the notation and previous definitions allowing to formulate the cosmic CHSH inequality.
@@
    To begin with, let $\alpha\in[0,\pi]$ be the angle between $\boldsymbol{\hat{x}}$ and $\boldsymbol{\hat{y}}$, that is $\cos\alpha\equiv{\boldsymbol{\hat{x}}}\cdot{\boldsymbol{\hat{y}}}$, and denote by $F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}}),\alpha\right>$ the mean value of the product $F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}})$ all over the direction pairs that form a given constant angle $\alpha$. Regarding these measurement pairs, the only to be considered are the ones present in the data sets of the measures actually performed. Further, we will also use the simplified notation $F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}}),\alpha\right>\equiv\left<\alpha\right>$.
@@
    Next we consider four different values for the constant angle $\alpha$, that is $\alpha_{i}$, $i=1,2,3,4$, and the corresponding four mean values $\alpha_{i}\right>$. Given these four mean values, there exist (see Ref. [10]) four unit vectors, $\boldsymbol{\hat{a}},\,\boldsymbol{\hat{a}^{\prime}},\,\boldsymbol{\hat{b}},\,\boldsymbol{\hat{b}^{\prime}}$, such that ${\boldsymbol{\hat{a}}}\cdot{\boldsymbol{\hat{b}}}=\cos\alpha_{1}$, ${\boldsymbol{\hat{a}}}\cdot{\boldsymbol{\hat{b}^{\prime}}}=\cos\alpha_{2}$, ${\boldsymbol{\hat{a}^{\prime}}}\cdot{\boldsymbol{\hat{b}}}=\cos\alpha_{3}$, and ${\boldsymbol{\hat{a}^{\prime}}}\cdot{\boldsymbol{\hat{b}^{\prime}}}=\cos\alpha_{4}$.
@@ 2.3 Proof of the cosmic CHSH inequality
Furthermore, the action of a generic rotation $R$ remains free so that the mean values, $F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}}),\alpha\right>$, in Eq. (2.1) are expressed $F({R\boldsymbol{\hat{x}}})F({R\boldsymbol{\hat{y}}})\right>$. With this notation, the inequality (2.1) can be written in this alternative way:
@@
where the dependence on the angles ($\alpha_{i}$) remains implicit.
Then, we can have either $F(R{\boldsymbol{\hat{b}}})=-F(R{\boldsymbol{\hat{b}^{\prime}}})$ or $F(R{\boldsymbol{\hat{b}}})=F(R{\boldsymbol{\hat{b}^{\prime}}})$. But, because of the assumed causality, the measurement results for the directions $R{\boldsymbol{\hat{b}}}$ and $R{\boldsymbol{\hat{b}^{\prime}}}$ do not depend on whether we measure $R{\boldsymbol{\hat{a}}}$ or $R{\boldsymbol{\hat{a}^{\prime}}}$. This means that in the two terms of the right hand side of inequality (2.2), we must put the same equality, $F(R{\boldsymbol{\hat{b}}})=-F(R{\boldsymbol{\hat{b}^{\prime}}})$ or $F(R{\boldsymbol{\hat{b}}})=F(R{\boldsymbol{\hat{b}^{\prime}}})$. In each of the two cases either the first term in inequality (2.2) cancels or the last term one cancels. Therefore, the maximum value of $\left|\left\langle C\right\rangle\right|$ is two. This completes the proof of inequality (2.2).
The expected values $F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}}),\alpha\right>\equiv\left<\alpha\right>$ are obtained through the expressions:
[…]
where $P\left({X,Y,\alpha}\right)$ stands for the probability of getting the result of $X\left({\pm 1}\right)$ for $F({\boldsymbol{\hat{x}}})$ and $Y\left({\pm 1}\right)$ for $F({\boldsymbol{\hat{y}}})$ (see the $F$ function definition in Subsection 2.1), for instance $P\left({+,+,\alpha}\right)$ represents the probability to obtain both temperatures of the two directions subtending the angle $\alpha$ greater than the mean temperature.
@@ 3 Computational and statistical analysis involved in the cosmic CHSH inequalities
This section contains a summary of the main tasks, procedures and considerations we undertake to collect the necessary ingredients of the inequalities (2.1), that is the $F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}}),\alpha\right>\equiv\left<\alpha\right>$ quantities defined in (2.3). Those quantities were obtained and analysed in a previous paper [10] using the COBE satellite data sets [33]. However, despite the importance of this first-generation satellite CMB measurement, two new generation satellites have improved enormously the aforementioned measures. Both satellites have significantly enhanced in terms of accuracy of the measured magnitudes as well as in terms of angular resolution. These circumstances deserve our attention and advocate a new recalculation and processing of the expected values. Once these tasks are undertaken we will review their implications. After COBE mission, in June 2001, the Wilkinson Microwave Anisotropy Probe (WMAP) NASA Explorer mission was launched [34]. While COBE satellite has an angular resolution of $7^{\circ}$ (FWHM – Full width at half maximum –), the WMAP one produced the first fine-resolution of 0.2 degree supplying detailed maps. Compared to COBE mission it has 45 times the sensitivity and 33 times the angular resolution [34]. Over the course of 9 years, WMAP team has produced five generation data sets named DR1, DR2, DR3, DR4 and DR5 corresponding to one, three, five, seven and nine-year data release, respectively. Finally, with a temperature resolution of one part in ${10}^{6}$ at an angular resolution of about 10 minutes of arc, the European Space Agency (ESA) Planck satellite [35] has provided the best accurate data sets. There are four public data releases labelled PR1, PR2, PR3 and PR4 relating to the released years 2013, 2015, 2018 and 2021, respectively. For an interesting comparison between both satellites (WMAP vs Planck) see [49].

### v2

# Violation of Bell inequalities from Cosmic Microwave Background data

@@ Abstract
In a recent paper [R. Dale. R. Lapiedra, and J. A. Morales-Lladosa, Phys. Rev. D 107, 023506 (2023)] a cosmic-like Clauser-Horne-Shimony-Holt (CHSH) inequality was proved for the temperature fluctuations $\delta T$ of the perturbed Cosmic Microwave Background (CMB), assuming local realism. This inequality can be tested from observational data. In fact, no violation for it has been found from the CMB data provided by the COBE satellite in different sky directions. This is a result which, being negative, is not conclusive in relation to the possible violation of the CHSH inequality in a cosmological context. So, the preceding analysis needs to be extended by using more precise current CMB data, like those provided by WMAP and Planck [^1] Based on observations obtained with Planck (http://www.esa.int/Planck), an ESA science mission with instruments and contributions directly funded by ESA Member States, NASA, and Canada.
We dedicate this work to the memory of our colleague and friend, Professor Diego Sáez Milán, whose insight and passion for science continue to inspire us.
satellites. This is a task of considerable importance but, technically, much more involved in data treatments. In the present, assuming again local realism behind measurements and observational data, this extension is accomplished. The result is that a sound violation of the CHSH inequality is found, which would mean the failure of the assumed local realism in accordance with the quantum origin of the primordial temperature fluctuations in the framework of standard inflation and the orthodox interpretation of the Quantum Mechanics.
Keywords CMBR experiments  $\cdot$ Statistical sampling techniques  $\cdot$ Inflation and CMBR theory  $\cdot$ Quantum cosmology
@@ 1 Introduction
Hence, it is usually assumed that a primordial quantum distribution of cosmic fluctuations dominated a very early era of the Universe but that CMB observational data respond to a classical statistic treatment obeying a stochastic distribution [22]. Diverse proposals exist to explain how such a quantum to classical transition takes place [22, 23, 24, 25], which are not free of criticisms [26]. Moreover, it is important to assess the degree of quantumness, if any, that a scalar inflationary field and the observed CMB anisotropies could encode [27, 28]. The transition from the firstborn quantum era to a classical or semi-classical one is an important topic which continues being addressed today, both from the theoretical and observational points of views [8, 29]. Thus, the proposed theoretical models go forward to an inasmuch more interesting, but very ‘entangled’ issue in Cosmology which claims for an experimental resolution.
@@
The present work does not include new theoretical insight or discussion about the transition from a quantum phase to a radiation dominant era, nor about how the primordial fluctuations were imprinted as CMB fluctuations, and its posterior evolution in an expanding universe. On the contrary, to face the issue we concentrate on the possibility to establish the true nature, if any, of the CMB anisotropies from their own observation. We propose a new bet to solve this ‘entangled’ cosmic situation: to cut the ‘Gordian knot’ with the sword of experiment.
@@
The present paper focuses on a variant of Bell’s inequality [30], known as the CHSH inequality (named after its authors, [31]), whose proof has been revisited in [32]. This inequality has been suitably mimicked and adapted to deal with CMB anisotropies data; then it was named ‘cosmic CHSH inequality’. In Ref. [10] no violation of this inequality was reported from the analysis of the COBE mission data [33]. But a concluding result about the inequality fulfilment deserves fuller treatment and demands a great amount of data like those provided by the WMAP and Planck measurement results [34, 35]. Now, in both cases, we find a sound violation of the Bell inequalities, which could be interpreted as a failure of local realism. The realism assumption posits that measurements can reveal an objective physical property intrinsic to the CMB, which exists prior to and independently of any observation. The locality assumption asserts that measurements conducted in two spacelike-separated regions at the decoupling time cannot exert any physical influence on each other at any subsequent cosmic instant.
@@
Laboratory experiments concerning the violation of Bell inequalities (see [11, 36] and references therein) show favour to a quantum mechanical interpretation. Nevertheless it seems to us that the controversy between both quantum interpretations continues open, particularly in Cosmology [37, 38]. Thus, the present paper could serve as a cosmic referee, potentially motivating the exploration of loopholes [39, 40, 41, 42, 43, 44] and alternative classical interpretations [45, 46], or ruling them out [47, 48], thereby providing new insights into the longstanding debate. This revision improves readability, precision, and fluency while maintaining a formal tone appropriate for scientific writing.
@@
The paper is organised as follows. First, Section 2 examines the fundamental hypothesis of local realism, which underpins the proof of Bell’s theorem, and reconciles it with measurements of the CMB. This section also revisits the “cosmic CHSH inequality”. Section 3 addresses the calculation of probabilities, expected values, and statistical methods applied to these cosmic inequalities. The computational procedures are summarised in Section 4. Subsequently, Sections 5 and 6 analyse observational data from the WMAP and Planck missions, demonstrating violations of the cosmic CHSH inequality in both cases. Section 7 provides a concise summary of the results obtained using degraded maps, including the confirmation of these findings with the original maps. Finally, Section 8 discusses the assumptions considered and the physical conclusions derived. Supplementary computations and technical results are presented in Appendices A and B.
@@ 2.1 Local realism assumption
[…]
By definition, realism implies the existence of hidden variables, $\lambda$, predating the measurements, such that the dichotomised CMB anisotropy values ($\pm 1$) are well-defined functions of $\lambda$ and the corresponding measurement directions. Explicitly,
@@
After each pair of measurements is completed, a new run is initiated, without requiring the subsequent pair of measurements to be conducted in the same initial state as the previous one. In a laboratory-based CHSH-like experiment, the system is actively prepared in the desired initial state before each new run. In the cosmic case, however, the invariance of the initial conditions prior to any measurement is not a crucial factor. This is because successive CMB measurement pairs are taken along randomly selected sky directions that maintain a constant angular separation. The only requirement is the assignment of a value, either +1+1 or 1-1, to each measurement.
[…]
@@
A few points should be considered regarding the terminology employed in both the standard CHSH setup and the cosmic variant explored in this work. In both contexts, the term ‘observer’ denotes any agent (human or instrumental) capable of performing measurements and extracting the corresponding information.
In a relativistic framework, a ‘local observer’ at an event $A$ is represented by a future-directed timelike unit vector $u_{A}$ at $A$. Hence, any conventional laboratory test of Bell inequalities involves two space-like separated events $A$ and $B$, with a pair of local observers $u_{A}$ and $u_{B}$ stationed at each event. Each observer carries out two measurements (e.g., spin or polarisation projections) along two different spacelike directions, defined respectively by the unit vectors $\boldsymbol{\hat{a}},\boldsymbol{\hat{a}^{\prime}}$ at $A$ and $\boldsymbol{\hat{b}},\boldsymbol{\hat{b}^{\prime}}$ at $B$.
In contrast, our proposed cosmic experiment involves a single effective observer — the satellite — performing CMB anisotropy measurements from the present epoch. However, the observed data correspond to the conditions at the time of decoupling, and thus are associated with hypothetical local observers at that earlier epoch. These observers are grouped in two pairs, $(u_{A},u_{B})$ and $(u_{A^{\prime}},u_{B^{\prime}})$, each linked to space-like separated events.
Crucially, our setup requires that all four event pairs — $AB$, $^{\prime}, ABA^{\prime}B$, and $A^{\prime}B^{\prime}$ — be space-like separated. During each observational run, four angular ‘windows’ on the sky remain simultaneously open from the satellite’s perspective: $\boldsymbol{\hat{a}}\boldsymbol{\hat{b}},\boldsymbol{\hat{a}}\boldsymbol{\hat{b}^{\prime}},\boldsymbol{\hat{a}^{\prime}}\boldsymbol{\hat{b}},\boldsymbol{\hat{a}^{\prime}}\boldsymbol{\hat{b}^{\prime}}$. Each window defines a pair of events, spatially separated at the surface of last scattering, aligned along the satellite’s line of sight via the corresponding unit vectors.
Despite the conceptual differences between the usual and the cosmic CHSH setups, their experimental implementations share a key structural element: in both cases, the CHSH inequality is constructed from the outcomes of four distinct measurements, namely ${\cal O}_{a}$, ${\cal O}_{a^{\prime}}$, ${\cal O}_{b}$, and ${\cal O}_{b^{\prime}}$, regardless of the specific nature of the physical observable $\cal O$} (which may be spin, polarization, or, in our case, temperature anisotropy). In the thought-experimental version of our cosmic scenario, each observer performs a single measurement—specifically, the sign of their local temperature fluctuation: $\delta T_{A}$, $\delta T_{A^{\prime}}$, $\delta T_{B}$, and $\delta T_{B^{\prime}}$. At the observational level, these measurements correspond to present-day data collected by the satellite observer and derived from redshifted CMB intensity. The cosmic CHSH inequality is thus constructed from these local outcomes, under the assumption of realism—i.e., the existence of hidden variables—as outlined earlier in subsection 2.3. In this way, the structure of the standard CHSH proof can be faithfully adapted to the cosmological setting, and any observed violation of the cosmic CHSH inequality can be interpreted as evidence for the breakdown of local realism within this framework.
@@ 2.2 Statment of the cosmic CHSH inequality
Let us summarise the notation and previous definitions allowing to formulate the cosmic CHSH inequality.
@@
[…]
@@
    Next we consider four different values for the constant angle $\alpha$, that is $\alpha_{i}$, $i=1,2,3,4$, and the corresponding four mean values $\left<\alpha_{i}\right$>. Given these four mean values, there exist (see Ref. [10]) four unit vectors, $\boldsymbol{\hat{a}},\,\boldsymbol{\hat{a}^{\prime}},\,\boldsymbol{\hat{b}},\,\boldsymbol{\hat{b}^{\prime}}$, such that ${\boldsymbol{\hat{a}}}\cdot{\boldsymbol{\hat{b}}}=\cos\alpha_{1}$, ${\boldsymbol{\hat{a}}}\cdot{\boldsymbol{\hat{b}^{\prime}}}=\cos\alpha_{2}$, ${\boldsymbol{\hat{a}^{\prime}}}\cdot{\boldsymbol{\hat{b}}}=\cos\alpha_{3}$, and ${\boldsymbol{\hat{a}^{\prime}}}\cdot{\boldsymbol{\hat{b}^{\prime}}}=\cos\alpha_{4}$.
@@ 2.3 Proof of the cosmic CHSH inequality
$\left.\begin{aligned} \left|\left\langle C\right\rangle\right|&\equiv\left|\left\langle F({\boldsymbol{\hat{a}}})F({\boldsymbol{\hat{b}}}),\alpha_{1}\right\rangle+\left\langle F({\boldsymbol{\hat{a}}})F({\boldsymbol{\hat{b}^{\prime}}}),\alpha_{2}\right\rangle\right.\\ &+\left.\left\langle F({\boldsymbol{\hat{a}^{\prime}}})F({\boldsymbol{\hat{b}}}),\alpha_{3}\right\rangle-\left\langle F({\boldsymbol{\hat{a}^{\prime}}})F({\boldsymbol{\hat{b}^{\prime}}}),\alpha_{4}\right\rangle\right|\leq 2.\end{aligned}\right$. (1)
For a given quad of angles, by a generic rotation we mean a rotation that (i) maps the directions onto a set of actually observed data on the CMB sky, and (ii) excludes rotations that merely permute the observed directions among themselves.
@@
$\left.\begin{aligned} \left|\left\langle C\right\rangle\right|&\equiv\left|\left\langle F(R{\boldsymbol{\hat{a}}})\left[F(R{\boldsymbol{\hat{b}}})+F(R{\boldsymbol{\hat{b}^{\prime}}})\right]\right\rangle\right.\\ &+\left.\left\langle F(R{\boldsymbol{\hat{a}^{\prime}}})\left[F(R{\boldsymbol{\hat{b}}})-F(R{\boldsymbol{\hat{b}^{\prime}}})\right]\right\rangle\right|\leq 2,\end{aligned}\right$. (2)
where the dependence on the angles ($\alpha_{i}$) remains implicit.
Then, we can have either $F(R{\boldsymbol{\hat{b}}})=-F(R{\boldsymbol{\hat{b}^{\prime}}})$ or $F(R{\boldsymbol{\hat{b}}})=F(R{\boldsymbol{\hat{b}^{\prime}}})$. But, because of the assumed causality, the measurement results for the directions $R{\boldsymbol{\hat{b}}}$ and $R{\boldsymbol{\hat{b}^{\prime}}}$ do not depend on whether we measure $R{\boldsymbol{\hat{a}}}$ or $R{\boldsymbol{\hat{a}^{\prime}}}$. This means that in the two terms of the right hand side of inequality (2), we must put the same equality, $F(R{\boldsymbol{\hat{b}}})=-F(R{\boldsymbol{\hat{b}^{\prime}}})$ or $F(R{\boldsymbol{\hat{b}}})=F(R{\boldsymbol{\hat{b}^{\prime}}})$. In each of the two cases either the first term in inequality (2) cancels or the last term one cancels. Therefore, the maximum value of $\left|\left\langle C\right\rangle\right|$ is two. This completes the proof of inequality (2).
The expected values $\left<F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}}),\alpha\right>\equiv\left<\alpha\right$> are obtained through the expressions:
$\left.\begin{aligned} \left\langle{F\left({\boldsymbol{\hat{x}}}\right)F\left({\boldsymbol{\hat{y}}}\right),\alpha}\right\rangle&=P\left({+,+,\alpha}\right)+P\left({-,-,\alpha}\right)-P\left({+,-,\alpha}\right)-P\left({-,+,\alpha}\right),\end{aligned}\right$. (3)
@@ 3 Computational and statistical analysis involved in the cosmic CHSH inequalities
This section contains a summary of the main tasks, procedures and considerations we undertake to collect the necessary ingredients of the inequalities (1), that is the $\left<F({\boldsymbol{\hat{x}}})F({\boldsymbol{\hat{y}}}),\alpha\right>\equiv\left<\alpha\right$> quantities defined in (3). Those quantities were obtained and analysed in a previous paper [10] using the COBE satellite data sets [33]. However, despite the importance of this first-generation satellite CMB measurement, two new generation satellites have improved enormously the aforementioned measures. Both satellites have significantly enhanced in terms of accuracy of the measured magnitudes as well as in terms of angular resolution. These circumstances deserve our attention and advocate a new recalculation and processing of the expected values. Once these tasks are undertaken we will review their implications. After COBE mission, in June 2001, the Wilkinson Microwave Anisotropy Probe (WMAP) NASA Explorer mission was launched [34]. While COBE satellite has an angular resolution of $7^{\circ}$ (FWHM – Full width at half maximum –), the WMAP one produced the first fine-resolution of 0.2 degree supplying detailed maps. Compared to COBE mission it has 45 times the sensitivity and 33 times the angular resolution [34]. Over the course of 9 years, WMAP team has produced five generation data sets named DR1, DR2, DR3, DR4 and DR5 corresponding to one, three, five, seven and nine-year data release, respectively. Finally, with a temperature resolution of one part in ${10}^{6}$ at an angular resolution of about 10 minutes of arc, the European Space Agency (ESA) Planck satellite [35] has provided the best accurate data sets. There are four public data releases labelled PR1, PR2, PR3 and PR4 relating to the released years 2013, 2015, 2018 and 2021, respectively. For an interesting comparison between both satellites (WMAP vs Planck) see [49].